Two tube lights are installed in a room. Probability of light A failing is 0.20 while the probability of light B failing is 0.15. What is the probability that both the lights are working on a given day? (final value is rounded off to 2 decimal places)
Answer by ikleyn(52781)

.

From the condition,


    - the probability that tube light A is working at a given day is  1 - 0.2 = 0.8,  and

    - the probability that tube light B is working at a given day is  1 - 0.15 = 0.85.


Therefore, the probability that both tube lights are working at a given day is the product  0.8*0.85 = 0.68.    ANSWER


It is the exact / precise value, so rounding is not required.